# Online Decision-Focused Learning

### v1

@@ 4.1 Online decision-focused learning on the simplex
This setting encompasses various decision-making scenarios such as portfolio selection li2014online or mixed strategy design in multiplayer games syrgkanis2015fast . In this case, we choose the negative entropy
@@
$\tilde{w}_{t,i}(\theta)=\frac{\exp(-\alpha_{t}^{-1}g_{i}(\theta,X_{t}))}{\sum_{k\in[d]}\exp(-\alpha_{t}^{-1}g_{k}(\theta,X_{t}))}\;.$ (8)
@@
$\nabla\tilde{w}_{t}(\theta)=-\frac{1}{\alpha_{t}}\mathopen{}\mathclose{{}\left[\mathrm{diag}[\tilde{w}_{t}(\theta)]-\tilde{w}_{t}(\theta)\tilde{w}_{t}(\theta)^{\mathsf{\scriptscriptstyle T}}}\right]\nabla_{\theta}g(\theta,X_{t})\;.$ (9)
The regularized minimizer (8) and its Jacobian (9) can readily be used in Algorithm 1. In this case, the regret (4) incurred by Algorithm 1 can be bounded as follows.
@@
Assume H 1, H 2, H 3, $\mathcal{W}=\mathcal{W}_{0}$, $\mathcal{R}=\mathcal{R}_{0}$ and that Algorithm 1 is instantiated with a non-increasing sequence $\mathopen{}\mathclose{{}\left(\eta_{t}}\right)_{t\in[T]}$. Then, denoting $P_{T}=\sum_{t=2}^{T}\mathopen{}\mathclose{{}\left\|\vartheta_{t}-\vartheta_{t-1}}\right\|$,
@@
$T^{-1}\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}}\right]$ $\leqslant\frac{D_{\Theta}^{2}+D_{\Theta}\mathbb{E}\mathopen{}\mathclose{{}\left[P_{T}}\right]}{2T\eta_{T}}+\sum_{t\in[T]}\frac{25D_{\mathcal{Z}}G\eta_{t}}{32T\alpha_{t}^{2}}$
$+2D_{\mathcal{Z}}\sum_{t\in[T]}\alpha_{t}[1+2\ln(d)C_{0}\{\ln(2\alpha_{t}^{-1})+(1-\beta)\ln^{2}(\alpha_{t}\ln d)\}]+\xi\;.$
The term $P_{T}$–which ultimately depends on the sequence $(X_{1},\bar{g}_{1}(X_{1})),\ldots,(X_{t},\bar{g}_{T}(X_{T}))$– controls the variation of the problem through time. Proposition 1 makes clear that keeping the dynamic regret (4) low becomes difficult when the revealed cost vectors vary drastically over time. Our guarantee does not depend on the dimension $m$ of $\Theta$, a notable strength stemming from the optimistic approach. This is sharp contrast with non-optimistic algorithms, such as DF-FTPL in Appendix B, which suffers a $m^{2}$ factor in its associated static regret bound (Proposition 4).
In addition to the usual trade-off associated with the learning rate $\eta_{t}$, Proposition 1 highlights a new trade-off introduced by the regularization strength $\alpha_{t}$. As $\alpha_{t}$ increases, the second term decreases as the mapping $\tilde{f}_{t}$ becomes smoother in the sense that its Lipschitz constant decreases (see Lemma 2). However, the third term grows larger as $\tilde{w}_{t}(\theta)$ becomes a poorer approximation of $w^{\star}_{t}(\theta)$. It is therefore natural to wonder how to pick the sequences $(\eta_{t})_{t\in[T]}$ and $(\alpha_{t})_{t\in[T]}$ to address these trade-offs.
In the following corollary, we define for any $t\in[T]$,
$P_{t}=\begin{cases}\sum_{s=1}^{t}\mathopen{}\mathclose{{}\left\|\vartheta_{s}-\vartheta_{s-1}}\right\|&\text{if }t\geqslant 2\;,\\ 0&\text{otherwise}\end{cases}\;.$
We prove that under a mild assumption regarding the growth of $(P_{t})_{t\in[T]}$, we can pick $(\alpha_{t})_{t\in[T]}$ and $(\eta_{t})_{t\in[T]}$ in a fully adaptative way to obtain an explicit regret bound.
Corollary 1.
Assume H 1,H 2, H 3, $\mathcal{W}=\mathcal{W}_{0}$, $\mathcal{R}=\mathcal{R}_{0}$ and $(t^{-1}(1+P_{t}))_{t\in[T]}$ is non-increasing almost surely. Then, using Algorithm 1 with $\alpha_{t}\propto t^{-1/4}(1+P_{t})^{1/4}$ and $\eta_{t}\propto t^{-3/4}(1+P_{t})^{3/4}$ for any $t\in[T]$ gives:
$\frac{\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}}\right]}{T}=\tilde{\mathcal{O}}\mathopen{}\mathclose{{}\left(\ln(d)\mathopen{}\mathclose{{}\left(\frac{1+\mathbb{E}\mathopen{}\mathclose{{}\left[P_{T}}\right]}{T}}\right)^{1/4}+\xi}\right)\;,$
where $\tilde{\mathcal{O}}$ contains polynomial dependencies in $\ln(1/\alpha),\ln(\ln(d))$.
Note that the resulting expected regret bound only depends logarithmically on the dimension of the decision space.
We stress that the assumption about $(t^{-1}(1+P_{t}))_{t\in[T]}$ being non-increasing can be weakened to the existence of a non-increasing sequence $(r_{t})_{t\in[T]}$ such that $t^{-1}P_{t}\leqslant r_{t}$ for any $t\in[T]$ without affecting the result. This assumption is merely technical, and we believe that Corollary 1 would hold without it up to slight modifications. Finally, under the setting of Corollary 1, we remark that $\eta_{t}$ and $\alpha_{t}$ are computable on the fly as they only require the knowledge of $P_{t}$, which depends itself on observed variables $(\vartheta_{s})_{s\in[T]}$. This differs, for instance, from the dynamic regret bound of OGD in (hazan2023introductiononlineconvexoptimization, , Chapter 10) which requires the knowledge of $P_{T}$ before training to reach their optimal rate.
We conclude by mentioning that in Appendix D, we consider a stronger version of H 2 which allows us to derive a counterpart of Proposition 1 and Corollary 1 for a stronger notion of dynamic regret.
4.2 Generic polytope
In Section 4.1, the very specific structure of the simplex $\mathcal{W}_{0}$ hinted us to choose $\mathcal{R}$ to be the negative entropy. However many problems, such as shortest-path, min-cost-flow lin2014quickest or unit commitment problems saravanan2013solution , require optimizing on a generic polytope. In what follows, we go back to the generic case where $\mathcal{W}=\mathcal{W}_{1}$ is a bounded, convex polytope with non-empty interior. We can write $\mathcal{W}_{1}$ as the intersection of $n$ half-spaces, denoting by $n>0$ its number of faces, that is
$\mathcal{W}_{1}=\{w\in\mathbb{R}^{d}:\>A^{\mathsf{\scriptscriptstyle T}}w-b\preccurlyeq 0\}\;,$
where $A\in\mathbb{R}^{d\times n},b\in\mathbb{R}^{n}$. In what follows, we assume that $\mathcal{W}_{1}$ is not degenerated, i.e. $AA^{\mathsf{\scriptscriptstyle T}}\in\mathbb{R}^{d\times d}$ is full-rank.
To tackle this more general structure, our strategy is to use a regularizer $\mathcal{R}$ such that $\tilde{w}_{t}$ always lies in the strict interior of $\mathcal{W}_{1}$, in which case the Jacobian map $\theta\mapsto\nabla\tilde{w}_{t}(\theta)$ can be determined by the first order condition and the implicit function theorem. To this end, we propose to use log-barriers as the regularizer in (5), that is
$\mathcal{R}_{1}:w\mapsto-\sum_{i=1}^{n}\ln(b_{i}-A_{i}^{\mathsf{\scriptscriptstyle T}}w)\;,$ (10)
where $A_{i}\in\mathbb{R}^{d}$ is the $i$-th column of $A$. With this choice of regularization, we show in Lemma 6 of Section F.2 that:
$\nabla\tilde{w}_{t}(\theta)=-\alpha_{t}^{-1}\mathopen{}\mathclose{{}\left(\sum_{i=1}^{n}(A_{i}^{\mathsf{\scriptscriptstyle T}}\tilde{w}_{t}(\theta)-b_{i})^{-2}A_{i}A_{i}^{\mathsf{\scriptscriptstyle T}}}\right)^{-1}\nabla_{\theta}g(\theta,X_{t})\;,$ (11)
which can be readily used in Algorithm 1. Note that (11) involves inverting a sum of rank-one matrices—a common operation in interior point methods—which can be performed efficiently; see jiang2020faster . With this choice, we can establish the following dynamic regret bound.
@@ 4.2 Generic polytope
Assume H 1, H 2, H 3, $\mathcal{W}=\mathcal{W}_{1}$, $AA^{\mathsf{\scriptscriptstyle T}}$ is full-rank, $\mathcal{R}=\mathcal{R}_{1}$ and Algorithm 1 is instantiated with a non-increasing sequence $\mathopen{}\mathclose{{}\left(\eta_{t}}\right)_{t\in[T]}$. Then,
$T^{-1}\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}}\right]$ $\leqslant\frac{D_{\Theta}^{2}+D_{\Theta}\mathbb{E}\mathopen{}\mathclose{{}\left[P_{T}}\right]}{2T\eta_{T}}+\sum_{t\in[T]}\frac{\eta_{t}}{2}\mathopen{}\mathclose{{}\left(\frac{GD_{\mathcal{Z}}\sup_{w\in\mathcal{W}_{1}}\mathopen{}\mathclose{{}\left\|b-Aw}\right\|_{\infty}^{2}}{\alpha_{t}\lambda_{\min}(AA^{\mathsf{\scriptscriptstyle T}})}}\right)^{2}$
$+2D_{\mathcal{Z}}\sum_{t\in[T]}\alpha_{t}[1+2n\max_{k\in[K]}\mathopen{}\mathclose{{}\left\|v_{k}}\right\|_{1}C_{0}(\ln(2\alpha_{t}^{-1})+(1-\beta)\ln^{2}(\alpha_{t}\ln d))]+\xi\;,$
where $P_{T}$ is defined as in Proposition 1.
Proposition 2 closely resembles to Proposition 1, but with an additional linear dependence on the number of faces $n>0$. This can be seen as the cost of operating over a more complex admissible space than the simplex. As before, it is possible to choose sequences $(\alpha_{t})_{t\in[T]}$ and $(\eta_{t})_{t\in[T]}$ to balance the trade-off discussed in Section 4.1 in a fully adaptative way.
Corollary 2.
Assume H 1, H 2, H 3, $\mathcal{W}=\mathcal{W}_{1}$, $\mathcal{R}=\mathcal{R}_{1}$ and $(t^{-1}(1+P_{t}))_{t\in[T]}$ is non-increasing almost surely. Then, using Algorithm 1 with $\alpha_{t}\propto t^{-1/4}(1+P_{t})^{1/4}$ and $\eta_{t}\propto t^{-3/4}(1+P_{t})^{3/4}$ for any $t\in[T]$ leads to :
$\frac{\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}}\right]}{T}=\tilde{\mathcal{O}}\mathopen{}\mathclose{{}\left(n\frac{1+\mathbb{E}\mathopen{}\mathclose{{}\left[P_{T}}\right]}{T}^{1/4}+\xi}\right)\;.$
Note that since we assume $AA^{\mathsf{\scriptscriptstyle T}}$ to be full rank, we have in particular that $n\geqslant d$, which is a deteriorated dependence compared to the $\ln(d)$ factor of Corollary 1.
5 Experiments
00111122223333444455550011112222333344445555Value of Item 1Value of Item 2
Figure 1: Figure 2 from mandi2024decision .
Setting.
@@ 5 Experiments
After having made their decision, the decision-maker observe the true item costs $\bar{g}_{t}(X_{t})=\mathbb{E}[v_{t}\,|\,X_{t}]$, and update $\theta_{t+1}$ for the next round based on this feedback. Note that this setting can directly be mapped in the simplex example of Section 4.2, since $a^{\star}_{t}(\theta_{t})=w^{\star}_{t}(\theta_{t})$ with $w^{\star}_{t}(\theta_{t})=\operatorname*{\arg\!\min}_{w\in\mathcal{W}_{0}}\mathopen{}\mathclose{{}\left\langle w,g(\theta_{t},X_{t})}\right\rangle$, $\mathcal{W}_{0}$ being the simplex of $\mathbb{R}^{K}$.
[…]
Synthetic data.
We instantiate the previous knapsack problem with the following synthetic data. For any $t\in[T]$, we draw $X_{t}\in\mathbb{R}^{K\times p}$ with correlated rows, and generate $v_{t}\in[0,1]^{k}$ as follows:
$v_{t}=\alpha\sin^{4}((2X_{t}\theta^{\star}_{t})^{-1})+\varepsilon_{t}\;,$ (12)
where $\alpha>0$, $\varepsilon_{t}\sim\mathrm{N}(0,I_{K})$ is a Gaussian noise and $\theta^{\star}_{t}\in\mathbb{R}^{p}$ is a parameter which follows an autoregressive recursion:
$\theta^{\star}_{t}=\frac{1}{2}\theta^{\star}_{t-1}+\frac{1}{2}\zeta_{t}\,,\quad\text{where}\quad\zeta_{t}\sim\mathrm{N}(0,I_{p})\;.$
This is a challenging data generating process, since $\bar{g}_{t}(X_{t})=\sin^{4}((2X_{t}^{\mathsf{\scriptscriptstyle T}}\theta^{\star}_{t})^{-1})$ is non-stationary and highly non-linear, and features are correlated. Equation 12 is discussed more in detail in Appendix C. To predict $\hat{v}_{t}$ from $X_{t}$, we assume that the decision-maker has access to a class of linear predictors of the form $g:(\theta,X)\mapsto X\theta$.
Comparison between DFL and PFL.
In this framework, we compare DF-OGD  (Algorithm 1) to a Prediction-Focused Online Gradient Descent (PF-OGD) described in Algorithm 2, which minimizes the loss $\ell^{\mathrm{mse}}:(\theta,v,X)\mapsto\mathopen{}\mathclose{{}\left\|v-X\theta}\right\|^{2}$.
@@
3:     Observe $X_{t}$, predict $\hat{v}_{t}=g(\theta_{t},X_{t})$
4:     Play $a^{\star}_{t}(\theta_{t})=\operatorname*{\arg\!\min}_{i\in[K]}\hat{v}_{t}(i)\;,$ observe $\bar{g}_{t}(X_{t})$.
5:     Update $\theta_{t+1}=\Pi_{\Theta}(\theta_{t}-\eta\nabla\ell^{\mathrm{mse}}(\theta_{t},\bar{g}_{t}(X_{t}),X_{t}))\;.$
@@
In Figure 2, we plot on the left-hand panel the average cumulated cost $t\mapsto t^{-1}\sum_{i=1}^{t}v_{t}(a^{\star}_{t}(\theta_{t}))$ incurred by DF-OGD and PF-OGD over 70 runs of $T=2000$ timesteps, as well as the associated 95% confidence intervals. On the right-hand panel, we plot the average Mean Square Error resulting from the sequence of prediction parameters $(\theta_{t})_{t\in[T]}$. Interestingly, DF-OGD produces parameters that induce better decisions as indicated by the lower average cumulated loss, but degraded predictions as suggested by the constantly higher MSE. This is in line with the intuition according to which the decision-focused approach sacrifices accuracy for the sake of better downstream decisions.
Figure 2: Average cumulated cost (left) and prediction errors (right) of DF-OGD and PF-OGD.
We conclude this section by mentioning that we perform additional numerical experiments to study how the relative performances of DFL and PFL are affected by model mispecification in Appendix C. This experiment empirically supports the idea that DFL outperforms PFL when models are clearly mispecified.
6 Conclusion.
Decision-focused learning is a promising approach for better leveraging prediction in real-life decision making processes. While the theoretical analysis of decision-focused learning is currently restricted to the batch setting, we extend it to the online framework. This allows us to consider non-stationary data processes and varying objective functions. We propose an algorithm based on regularization, perturbation and optimism which only requires picking a regularizer along with a sequence of regularization coefficients and learning rates. In particular, we show how to pick these hyper-parameters in two canonical setups, namely when the decision space is the simplex, and when it is a general bounded, convex polytope with non-empty interior. In both cases, we provide theoretical guarantees for our algorithm, in the form of upper bounds on the expected average dynamic regret. Finally, we compare the empirical performances of our approach to a more standard prediction-focused learning procedure in a knapsack experiment with a mispecified prediction model. In this experiment, our method outperforms the PFL one, which supports the idea that DFL leads to decision that are more robust to model mispecification.
We believe that this work can be extended in several ways. First, it would be interesting to study whether meaningful guarantees could be obtained under weaker assumptions than ours. Second, we believe that alternative smoothing techniques could be successfully employed, such as Moreau-Yosida transform and proximal operator, or random perturbation and Fenchel conjugate techniques berthet2020learning . Analyzing the theoretical performance of these methods would be of significant interest. Third, it would be valuable to investigate less adversarial environments, such as those involving i.i.d. data, to explore whether stronger theoretical guarantees could be obtained and whether novel algorithmic designs might emerge. Finally, implementing our method in more ambitious experimental settings would be of great interest from a practical point of view.
Appendix A A brief recap on online learning
Online learning (OL) is a paradigm in which data is processed sequentially, either because the practitionner does not collect all data prior to analysis or because the dataset dynamically evolves through time, or simply because handling batch of data is numerically too demanding. From the seminal work of [77], which proposed an online version of the celebrated gradient descent algorithm, OL has been intensively studied (see [36] for an overview). The classical performance criterion of an online learning algorithm is the static regret. Given a sequence of loss functions $(\ell_{t}:\mathcal{K}\rightarrow\mathbb{R})_{t\geqslant 1}$, the static regret compares the efficiency of a sequence of decisions $\hat{\mu}=(\hat{\mu}_{t})_{t\geqslant 1}$ to the best fixed strategy for $T>0$:
$\operatorname{Regret}_{T}^{\mathrm{S}}(\hat{\mu})=\sum_{t=1}^{T}\ell_{t}(\hat{\mu}_{t})-\inf_{\mu_{0}\in\mathcal{K}}\sum_{t=1}^{T}\ell_{t}(\mu_{0}),$ (13)
Classical upper bounds on static regret involve a sub-linear rate. For instance, [77] proposed a $\mathcal{O}(\sqrt{T})$ bound for Online Gradient Descent (OGD) which is valid for convex losses.
Static regret may not be sufficient to assert the efficiency of an online algorithm as the class of static strategies is limited compared to all possible strategies. Hence the notion of dynamic regret introduced by [77] and further developed by [34]. For any sequence $\hat{\mu}$ of predictors and any sequence $\mu$ of dynamic strategies, the dynamic regret is given by
$\operatorname{Regret}_{T}(\hat{\mu},\mu)=\sum_{t=1}^{T}\ell_{t}(\hat{\mu}_{t})-\sum_{t=1}^{T}\ell_{t}(\mu_{t})\,,\qquad T\geq 1\,.$
Dynamic regret has attracted a lot of interest recently, especially when the comparator sequence is set to $(\inf_{\mu\in\mathcal{K}}\ell_{t}(\mu))_{t\geqslant 1}$ (worst-case dynamic regret, as in [13, 74]) but also for any comparator sequence (universal dynamic regret, as in [75]). Note that dynamic regret is a stronger notion that static regret, since it compares the performance of $\hat{\mu}$ to the best dynamic strategy ($\inf_{\mu\in\mathcal{K}}\ell_{t}(\mu)$), while static regret relies on the best static strategy $(\inf_{\mu\in\mathcal{K}}\sum_{t=1}^{T}\ell_{t}(\mu))$.
Appendix B An online decision-focused algorithm with static regret guarantees
In our work, we mainly focused on upper bounds of the dynamic regret, given its versatility and its ability to accommodate non-stationary environments. However, static regret as defined in (13) is a more common measure of performance. In this subsection, we show that it is possible to extend our analysis to static regret based on techniques from [64]. More precisely, we propose in what follows a decision-focused version of the Follow-The-Perturbed-Leader (FTPL) algorithm for nonconvex losses. We first introduce some necessary background in Section B.1, and apply it to Decision-Focused learning in Section B.2.
B.1 Background on FTPL.
Approximate Optimization Oracle.
We denote the output of such an optimization oracle by $\mu\star=\mathcal{O}_{\xi,\chi}(\ell-\langle\sigma,\cdot\rangle)$. Note the property (14) is very close from the assumption H 3 we made in Section 3 on the expert sequence $(\vartheta_{t})_{t\in[T]}$ (in this case $\chi=0$ as no $\sigma$ is involved).
Follow The Perturbed Leader.
where $\sigma_{t}\in\mathbb{R}^{d}$ is a random perturbation such that $\sigma_{t,j}$, the j-th ccoordinate of $\sigma_{t}$, is sampled from $\operatorname{Exp}(\eta)$, the exponential distribution with parameter $\eta$.
Proposition 3 (Theorem 1 of [64]).
Let $D$ be the diameter of $\mathcal{K}$. Suppose that $\ell_{t}$ is $L$-Lipschitz w.r.t L1 norm, for all $t\in[T]$. For any fixed $\eta$, FTPL (Equation 15) with access to a $(\xi,\chi)$-approximate" optimization oracle satisfies the following static regret bound:
$\frac{1}{T}\operatorname{Regret}_{T}^{\mathrm{S}}(\hat{\mu})\leqslant\mathcal{O}\mathopen{}\mathclose{{}\left(\eta m^{2}DL^{2}+\frac{m(\chi T+D)}{\eta T}+\xi+\chi mL}\right),$
where $\operatorname{Regret}_{T}^{\mathrm{S}}(\hat{\mu})$ is defined as in (13), and $\mathbb{E}$ denotes the expectation over $\sigma_{1}\cdots\sigma_{T}$.
B.2 Application to Decision-Focused Learning.
We now show how to adapt the techniques and results from Section B.1 to our online decision-focused problem. First, we instantiate the static regret defined (13) in our framework. For any $T>0$, loss sequence $(f_{t})_{t\in[T]}$ as defined in (3), and algorithms $(\mathsf{Alg}_{t})_{t\in[T]}$ outputting $(\theta_{t})_{t\in[T]}$, we define for any $\theta\in\Theta$:
$\mathfrak{R}_{T}^{\mathrm{S}}(\theta)=\sum_{t\in[T]}F_{t}(\theta_{t})-\sum_{t\in[T]}F_{t}(\theta)\;.$ (16)
where $F_{T}:\theta\rightarrow\mathbb{E}\mathopen{}\mathclose{{}\left[f_{t}(\theta)\mid\mathcal{H}_{t-1}}\right]$ is the conditional expectation of the loss $f_{t}$ with respect to the history $\mathcal{H}_{t-1}$.
Note that Proposition 3 holds for any sequence of lipschitz losses, and that the surrogate losses ($\tilde{f}_{t}$) are indeed Lipschitz for our choice of regularizers in Section 4.1 and Section 4.2 (see Lemmas 2 and 7). from this observation, we derive a non-optimistic online algorithm following the same pattern as DF-OGD (Algorithm 1), but which leverages FTPL to enjoy the guarantee proposition 3.
Algorithm 3 Decision-Focused Follow The Perturbed Leader DF-FTPL

### v2

@@ C.1 Additional framework
This setting encompasses various decision-making scenarios such as portfolio selection [47] or mixed strategy design in multiplayer games [68]. In this case, we choose the negative entropy
@@
$\tilde{w}_{t,i}(\theta)=\frac{\exp(-\alpha^{-1}g_{i}(\theta,X_{t}))}{\sum_{k\in[d]}\exp(-\alpha^{-1}g_{k}(\theta,X_{t}))}\;.$
@@
$\nabla\tilde{w}_{t}(\theta)=-\frac{1}{\alpha}\mathopen{}\mathclose{{\left[\mathrm{diag}[\tilde{w}_{t}(\theta)]-\tilde{w}_{t}(\theta)\tilde{w}_{t}(\theta)^{\mathsf{\scriptscriptstyle T}}}}\right]\nabla_{\theta}g(\theta,X_{t})\;.$ (12)
C.2 Results
DF-FTPL
@@ C.2 Results
Assume H˜1, H˜2 and having access to an $\xi$-approximate optimization oracle $\mathbf{O}_{\xi}$ adapted to $\mathopen{}\mathclose{{\left\{\sum_{i=1}^{t}\tilde{f}_{i}-\mathopen{}\mathclose{{\left\langle\sigma_{t},\cdot}}\right\rangle}}\right\}_{t\in[T]}$. Fix $\mathcal{W}=\mathcal{W}_{0}$, $\mathcal{R}=\mathcal{R}_{0}$. Let $\{\theta_{t}\}_{t\in[T]}$ be the output of DF-FTPL (Algorithm˜1) instantiated with learning step $\eta>0$ and regularization coefficients $\alpha_{t}=\alpha>0$ for any $t$. Then:
@@
$T^{-1}\mathbb{E}\mathopen{}\mathclose{{\left[\mathfrak{R}^{s}_{T}}}\right]$ $\leqslant\tilde{\mathcal{O}}\mathopen{}\mathclose{{\left(\eta m^{2}D\frac{1}{\alpha^{2}}+\frac{mD}{\eta T}+\xi+\alpha\ln(d)}}\right)\;,$
The proof of this result is postponed to Section˜G.4.
DF-OGD
@@
Assume H˜1, H˜2, and having access to an $\xi$-approximate optimization oracle $\mathbf{O}_{\xi}$ adapted to $\mathopen{}\mathclose{{\left\{\tilde{f}_{i}}}\right\}_{t\in[T]}$. Fix $\mathcal{W}=\mathcal{W}_{0}$, $\mathcal{R}=\mathcal{R}_{0}$. Let $\{\theta_{t}\}_{t\in[T]}$ be the output of DF-OGD (Algorithm˜2) instantiated with the non-increasing sequence $\mathopen{}\mathclose{{\left(\eta_{t}}}\right)_{t\in[T]}$ and regularization coefficients $(\alpha_{t})_{t\in[T]}$. Then:
$T^{-1}\mathbb{E}\mathopen{}\mathclose{{\left[\mathfrak{R}^{d}_{T}}}\right]$ $\leqslant\frac{D_{\Theta}^{2}+D_{\Theta}\mathbb{E}\mathopen{}\mathclose{{\left[P_{T}}}\right]}{2T\eta_{T}}+\sum_{t\in[T]}\frac{25D_{\mathcal{Z}}G\eta_{t}}{32T\alpha_{t}^{2}}$
where $P_{T}=\sum_{t=1}^{T-1}\mathopen{}\mathclose{{\left\|\vartheta_{t+1}-\vartheta_{t}}}\right\|$.
The proof of this result is postponed to Section˜G.5.
Note that the principal gain, compared to the general polytope case, is to avoid the dependency in $n$ the number of faces and enjoys a nice $\ln(d)$ dependency in the dimension of the decision space.
Appendix D Background on the FTPL algorithm
Approximate Optimization Oracle.
We denote the output of such an optimization oracle by $\mu^{*}=\mathcal{O}_{\xi,\chi}(\ell-\langle\sigma,\cdot\rangle)$. Note the notion of oracle described in (13) is very close from the Definition˜1 we made in Section˜3 on the expert sequence $(\vartheta_{t})_{t\in[T]}$ (in this case $\chi=0$ as no $\sigma$ is involved).
Remark 2.
Note that, in our work, we made the choice to fix $\chi=0$. This stronger assumption is due to the will of having a unifying framework encompassing both the setup for DF-FTPL and DF-OGD.
Follow The Perturbed Leader.
where $\sigma_{t}\in\mathbb{R}^{d}$ is a random perturbation such that $\sigma_{t,j}$, the j-th coordinate of $\sigma_{t}$, is sampled from $\operatorname{Exp}(\eta)$, the exponential distribution with parameter $\eta$.
Proposition 4 (Theorem 1 of [67]).
Let $D$ be the diameter of $\mathcal{K}$. Suppose that $\ell_{t}$ is $L$-Lipschitz w.r.t L1 norm, for all $t\in[T]$. For any fixed $\eta$, FTPL (Equation˜14) with access to a $(\xi,\chi)$-approximate" optimization oracle satisfies the following static regret bound:
$\frac{1}{T}\sum_{t=1}^{T}\ell_{t}(\hat{\mu}_{t})-\inf_{\mu}\sum_{t=1}^{T}\ell_{t}(\mu)\leqslant\mathcal{O}\mathopen{}\mathclose{{\left(\eta m^{2}DL^{2}+\frac{m(\chi T+D)}{\eta T}+\xi+\chi mL}}\right),$
where $\mathbb{E}$ denotes the expectation over $\sigma_{1}\cdots\sigma_{T}$.
Appendix E Additional experimental material
The motivating example from [54].
@@ Appendix E Additional experimental material
0112233445501122334455Value of Item 1Value of Item 2
Figure 2: Figure 2 from [54].
Details on the experimental setup.
In this paragraph, we provide more detail about the experimental setup in Section˜4 and perform additional numerical simulations. We start by explaining more precisely how the data used in our experiment are drawn. First, for any $t\in[T]$, $X_{t}\in\mathbb{R}^{K\times p}$ is constructed as follows: (i) we generate a Toeplitz covariance matrix $\Sigma=(\rho^{\mid i-j\mid})_{(i,j)\in[K]^{2}}$ for some $\rho\in(0,1)$, (ii) apply a Cholesky decomposition to it: $\Sigma=LL^{\mathsf{\scriptscriptstyle T}}$, and (iii) generate a matrix $\bar{X}_{t}$ with standard Gaussian entries. Then, $X_{t}$ is defined as $X_{t}=L\bar{X}_{t}$. This introduces correlation between features, which makes convergence harder for an online linear model. Second, we generate $v_{t}\in\mathbb{R}^{K}$ as follows:
$c_{t}=\min\mathopen{}\mathclose{{\left(\max\mathopen{}\mathclose{{\left(\tilde{c}_{t},0}}\right),1}}\right)\quad\text{with}\quad\tilde{c}_{t}=A\sin^{4}((2X_{t}\theta^{\star}_{t})^{-1})+\varepsilon\;.$
In the above equation, $A>0$ is a constant, $\theta_{t}^{\star}$ satisfies:
$\theta^{\star}_{t}=\frac{1}{2}\theta^{\star}+\frac{1}{2}\zeta_{t}\,,\quad\text{where}\quad\zeta_{t}\sim\mathrm{N}(0,I_{K})\quad\text{and}\quad\theta^{\star}\in\mathbb{R}^{K}\;,$
We now present the values used for the different parameters in our experiment. We consider $K=5$ items and $p=10$ features. The horizon is set to $T=5000$. For each plot, we run $N=10$ times DF-OGD and PF-OGD. The reported error bars are 95% Gaussian confidence intervals (from the CLT). In (9), $A=45$ and the correlation coefficient for $X_{t}$ is $\rho=0.8$. Our algorithms are instantiated with the following parameters. First, DF-OGD runs with $(\alpha_{t})$ and $(\eta_{t})$ as suggested by our theoretical analysis. The oracle from H˜2 is obtained through a SGD algorithm performing $t_{\vartheta}=10$ steps at each iteration with a learning rate set to $\eta_{\vartheta}=0.01$, while being initialized at $\vartheta_{t-1}$. On the other hand, PF-OGD runs with a learning rate $\eta=10$ determined by grid search. Our code is implemented with pytorch.
Benchmark algorithms.
@@
3:  Observe $X_{t}$, predict $c_{t}=g(\theta_{t},X_{t})$
4:  Play $v^{\star}_{t}(\theta_{t})=\operatorname*{\arg\!\min}_{i\in[K]}c_{t,i}\;,$ observe $\bar{g}_{t}(X_{t})$.
5:  Update $\theta_{t+1}=\Pi_{\Theta}(\theta_{t}-\eta\nabla\ell^{\mathrm{mse}}(\bar{g}_{t}(X_{t}),X_{t}\theta_{t}))\;.$
@@
Deviation from realizability.
As argued in Section˜1, prediction-focused learning would be optimal if the prediction model made no mistake, since in this case problem (2) would perfectly align with (1). As a consequence, one would expect prediction-focused learning to perform very well in the realizable setting, that is when for any $t\in[T]$, there exists $\bar{\theta}_{t}\in\Theta$ such that $\bar{g}_{t}(X_{t})=g(\bar{\theta}_{t},X_{t})$. On the flip side, it is likely to struggle in a mispecified setting. This is in contrast with DF-OGD, whose theoretical performance analysis does not highlight any peculiar dependence on realizability. We therefore hypothetize that PFL should outperform DFL in the realizable case, and DFL should take the upper hand when the prediction model is highly mispecified. To test this hypothesis, we run an alternative experiment where we interpolate between the well-specified, linear case, and ill-specified, non-linear one. More precisely, for $\gamma\in[0,1]$, we generate $(X_{t},v^{(\gamma)}_{t})_{t\in[T]}$ as follows:
In Figure˜3, we make $\gamma$ vary from 0 to 1. At each value, we plot the average cumulated cost gap between DF-OGD and PF-OGD, averaged over $50$ runs with horizon $T=1000$. This experiment provides empirical support to the previous conjecture, namely, DFL becomes a competitive option when the prediction model is ill-specified.
Figure 3: Average cumulated cost gap PFL - DFL as a function of $\gamma$.
Algorithm 4 Online Smart Predict-then-Optimize Online Gradient Descent (Online SPO).